Grade 11Physical Chemistry

A solution is prepared by adding 6g of a solute 'X' in 120g of water calculate the mass percentage of the solute.

Aruna
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
mass of solute = 6 gm
mass of solution = 120 + 6 = 126
 
hence mass percentage of solute = (6/126) * 100 = 100/21 = 4.762 %